OpenStudy (anonymous):

During a race, a runner runs at a speed of 6 m/s. 2 seconds later, she is running at a speed of 10 m/s. What is the runner’s acceleration. also can you show ur work so i understand

OpenStudy (anonymous):

You are given\[v_i = 6\frac{ \text{m} }{ \text{s} }\]\[v_f = 10\frac{ \text{m} }{ \text{s} }\]\[\Delta t = 2\text{s}\]Acceleration is calculated using\[a = \frac{ \Delta v }{ \Delta t } = \frac{ v_f - v_i }{ \Delta t }\]Substitute in the values and solve for \(a\)